Fickel Company has two manufacturing departments-Assembly and Testing & Packaging. The predetermined overhead rates in Assembly and Testing & Packaging are $16.00 per direct labor-hour and $12.00 per direct labor-hour, respectively. The companys direct labor wage rate is $20.00 per hour. The following information pertains to Job N-60: Testing & Packaging Assembly 340 25 Direct materials 180 Direct labor 40 Required: 1. What is the total manufacturing cost assigned to Job N-60? 2. If Job N-60 consists of 10 units, what is the unit product cost for this job? (Round your answer to 2 decimal places.) 1. Total manufacturing cost 2. Unit product cost per unit
Answer #1
Concepts and reason

Manufacturing Cost: It is the sum of all the cost related to the production of a product. It comprises of Variable Cost, Fixed Cost, Manufacturing Cost and Mixed Cost, etc. It is calculated by sum of all the Cost related to the product.

Unit cost: It can be referred as the costing of the product manufactured on an individual product basis. The total cost incurred on the manufacturing of the product is divided by the number of units produced and the resultant figure is the unit cost of the product.

Fundamentals

Direct Labour: Workers or employees directly involved in the manufacture of goods and services are called the direct labour. Direct labour can be directly traceable to a specific product or cost centre. In other words, the cost of the workers which is directly involved in the transformation of the raw material into the finished product.

Direct Materials: These are the materials which are used in the Production process which constitutes an integral portion of the final product, the cost of which is identifiable and traceable directly to it. The cost of direct material can be directly chargeable to the final product as compared to the indirect material.

Overhead Cost: It can be referred as an accounting term which states all the indirect expenses incurred by an organization for operating a business. These expenses are not directly related to the manufacturing of a product or rendering of a service.

Predetermined Overhead Rate: It is that rate based on which the overhead expenses are absorbed over the period. It is generally a blanket rate which is further multiplied with the most important factor of production which can be labour hours or machine hours to calculate the overhead expense for the year.

1.

Compute the overheads allocated to assembly department, using the equation as shown below:

Allocatedoverheads=LaborcostWagerate×Predeterminedrate=$180$20perhour×$16perhour=$144\begin{array}{c}\\{\rm{Allocated overheads}} = \frac{{{\rm{Labor cost}}}}{{{\rm{Wage rate}}}} \times {\rm{Predetermined rate}}\\\\ = \frac{{{\rm{\$ 180}}}}{{{\rm{\$ 20 per hour}}}} \times {\rm{\$ 16 per hour}}\\\\ = {\rm{\$ 144}}\\\end{array}

Hence, the overheads allocated to assembly department is $144.

Compute the overheads allocated to testing and packing department, using the equation as shown below:

Allocatedoverheads=LaborcostWagerate×Predeterminedrate=$40$20perhour×$12perhour=$24\begin{array}{c}\\{\rm{Allocated overheads}} = \frac{{{\rm{Labor cost}}}}{{{\rm{Wage rate}}}} \times {\rm{Predetermined rate}}\\\\ = \frac{{{\rm{\$ 40}}}}{{{\rm{\$ 20 per hour}}}} \times {\rm{\$ 12 per hour}}\\\\ = {\rm{\$ 24}}\\\end{array}

Hence, the overheads allocated to testing and packing department is $24.

Compute the total overhead cost applied to Job N-60, using the equation as shown below:

Totaloverheads=Assemblyoverheads+Testingoverheads=$144+$24=$168\begin{array}{c}\\{\rm{Total overheads}} = {\rm{Assembly overheads}} + {\rm{Testing overheads}}\\\\ = {\rm{\$ 144}} + {\rm{\$ 24}}\\\\ = {\rm{\$ 168}}\\\end{array}

Hence, the total overheads cost applied to Job N-60 is $168.

Compute the total direct material cost applied to Job N-60, using the equation as shown below:

Totalmaterials=Assemblymaterials+Testingmaterials=$340+$25=$365\begin{array}{c}\\{\rm{Total materials}} = {\rm{Assembly materials}} + {\rm{Testing materials}}\\\\ = {\rm{\$ 340}} + {\rm{\$ 25}}\\\\ = {\rm{\$ 365}}\\\end{array}

Hence, the total material cost applied to Job N-60 is $365.

Compute the total direct labor cost applied to Job N-60, using the equation as shown below:

Totallabor=Assemblylabor+Testinglabor=$180+$40=$220\begin{array}{c}\\{\rm{Total labor}} = {\rm{Assembly labor}} + {\rm{Testing labor}}\\\\ = {\rm{\$ 180}} + {\rm{\$ 40}}\\\\ = {\rm{\$ 220}}\\\end{array}

Hence, the total labor cost applied to Job N-60 is $220.

Compute the total manufacturing cost applied to Job N-60, using the equation as shown below:

Totalmanufacturingcost=Totalmaterialcost+Totallaborcost+Totaloverheads=$365+$220+$168=$753\begin{array}{c}\\{\rm{Total manufacturing cost}} = {\rm{Total material cost}} + {\rm{Total labor cost}} + {\rm{Total overheads}}\\\\ = {\rm{\$ 365}} + {\rm{\$ 220}} + {\rm{\$ 168}}\\\\ = {\rm{\$ 753}}\\\end{array}

Hence, the total manufacturing cost applied to Job N-60 is $753.

2.

Compute the unit product cost, using the equation as shown below:

Unitcost=TotalmanufacturingcostUnitsproduced=$75310units=$75.30\begin{array}{c}\\{\rm{Unit cost}} = \frac{{{\rm{Total manufacturing cost}}}}{{{\rm{Units produced}}}}\\\\ = \frac{{{\rm{\$ 753}}}}{{{\rm{10 units}}}}\\\\ = {\rm{\$ 75}}{\rm{.30}}\\\end{array}

Ans: Part 1

The total manufacturing cost applied to Job N-60 is $753.

Part 2

The unit product cost is $75.30.